OpenStudy (anonymous):

The ages of the eight-member school board are listed below. 28, 40, 34, 46, 60, 34, 51, 57 If A is equal to the median age of the group and B is equal to the mode of the group/s age, what is the value of A-B?

OpenStudy (anonymous):

Median is 43 Mode is 34

OpenStudy (anonymous):

The ages of the eight-member school board are listed below. 28, 40, 34, 46, 60, 34, 51, 57 If A is equal to the median age of the group and B is equal to the mode of the group/s age, what is the value of A-B? 6 7 8 9 10

OpenStudy (anonymous):

A - B (subtract B from A) at least that much u can do!!!!!!!!!!

OpenStudy (anonymous):

idk bruh.... lol

OpenStudy (anonymous):

im flutterin around wit yu

OpenStudy (anonymous):

flutterin*

OpenStudy (anonymous):

43 - 34 = 9..................who u r

OpenStudy (anonymous):

1, 8,15, 22, . . . What number is next in the above sequence? 37 30 29 27 none of the above

OpenStudy (anonymous):

29

OpenStudy (anonymous):

i doubt it may be 30

OpenStudy (anonymous):

hmmmm may be 27

OpenStudy (anonymous):

f 7a = (91)(13), what is the value of \[\sqrt{a}\] ? Grid-in type answer.

OpenStudy (anonymous):

13

OpenStudy (anonymous):

\[2x ^{3}y ^{4}z ^{3}a ^{2}\over4x ^{2}y ^{5}za\]

OpenStudy (anonymous):

may i know asl plz

OpenStudy (anonymous):

18 m fl

OpenStudy (anonymous):

kin you please answer the last question

OpenStudy (anonymous):

\[\LARGE \begin{array}{l} \left( {{{\left( {\frac{x}{7}} \right)}^2}\sqrt {\frac{{\left| {|x| - 3} \right|}}{{|x| - 3}}} + {{\left( {\frac{y}{3}} \right)}^2}\sqrt {\frac{{\left| {y + \frac{{3\sqrt {33} }}{7}} \right|}}{{y + \frac{{3\sqrt {33} }}{7}}}} - 1} \right) - \left( {\left| {\frac{x}{2}} \right| - \left( {\frac{{3\sqrt {33} - 7}}{{112}}} \right){x^2} - 3 + \sqrt {1 - {{(\left| {\left| x \right| - 2} \right| - 1)}^2}} - y} \right) \cdot \\ \left( {9\sqrt {\frac{{\left| {(\left| x \right| - 1)(\left| x \right| - 0.75)} \right|}}{{(1 - \left| x \right|)(\left| x \right| - 0.75)}}} - 8\left| x \right| - y} \right) \cdot \left( {3\left| x \right| + 0.75\sqrt {\frac{{\left| {(\left| x \right| - 0.75)(\left| x \right| - 0.5)} \right|}}{{(0.75 - \left| x \right|)(\left| x \right| - 0.5)}}} - 8\left| x \right| - y} \right) \cdot \\ \left( {2.25\sqrt {\frac{{\left| {(x - 0.5)(x + 0.5)} \right|}}{{(0.5 - x)(x + 0.5)}}} - y} \right) \cdot \left( {\frac{{6\sqrt {10} }}{7} + (1.5 - 0.5\left| x \right|)\sqrt {\frac{{\left| {\left| x \right| - 1} \right|}}{{\left| x \right| - 1}}} - \frac{{6\sqrt {10} }}{{14}}\sqrt {4 - {{(\left| x \right| - 1)}^2}} - y} \right) = 0 \end{array}\]